# Representation Learning and Identity Adversarial Training for Facial Behavior Understanding

### v1

@@ 5 Analysis of Identity Adversarial Training
In this section, we elucidate IAT by first showing the identity learning issue in AU tasks (Sec. 5.1), then demonstrating the learning dynamics refined by IAT (Sec. 5.2), and finally illustrating the importance of ensuring a strong regularization of IAT (Sec. 5.3).
@@ 5.1 Linear probing for identity recognition
To demonstrate the identity learning phenomenon in AU detection, we apply linear probing [14] on a trained AU detection model (FMAE) and evaluate the identity recognition accuracy on the BP4D dataset, which contains 41 subjects. Specifically, we freeze the backbone $G_{f}(\cdot;\theta_{f})$ of a well-trained FMAE and add a learnable linear classifier on top of the backbone to predict the identity label. For each subject in BP4D, we randomly draw 70 samples for training and 30 samples for testing. The resultant accuracy under linear probing is shown in Fig. 3, the red line indicates that FMAE can recognize more than half of people among the 41 subjects even though the model is only trained for one epoch. Given enough training, the identity recognition accuracy can be as high as 83%. By contrast, IAT significantly alleviates this identity learning issue with 4.6% accuracy after one epoch of training and 27.9% accuracy at epoch 19.
Figure 3: Identity accuracy (%) evaluated by linear probing on the BP4D dataset. IAT greatly reduces the identity-related features learned by the network backbone $G_{f}(\cdot;\theta_{f})$.
An interesting phenomenon is that even under the strong identity unlearning regularization, FMAE-IAT seems still to partially learn the identity-based features, by showing 27.9% accuracy (higher than the random guess accuracy 2.4%). We believe that the inherent high correlation between training and testing images for each subject provides the possibility for the model to infer the identity by looking at the non-face area.
We also visualize the feature output from the backbone of FMAE and FMAE-IAT using t-SNE [68] and see how these features are clustered according to the identity label. Fig. 4 presents the t-SNE results for 20 subjects in BP4D (41 subjects in total), given trained FMAE and FMAE-IAT models. It is clear that the identity-based feature clusters in FMAE become less linearly distinguishable (the ID head is a linear layer) under the effect of IAT.
@@
5.2 F1 learning dynamics
After showing that a regular AU model (FMAE) learns the subject identity, we now illustrate that the identity shortcut learning leads to a trivial AU prediction solution that is inferior to the solution delivered by IAT. Concretely, we observe that FMAE and FMAE-IAT have totally different learning dynamics in terms of AU predictions (indicated by F1 score). Fig. 5 shows the F1 score of both models along the training epochs, where the two models share the same learning rate, batch size and initial training states. It is clear from Fig. 5 that FMAE is optimized quickly and converges at the third epoch with an F1 score of 65.45% under the identity shortcut. In contrast, FMAE-IAT learns the AU decision boundary progressively and converges only at epoch 15 with an F1 score of 66.66%. One can infer that IAT explicitly pushes the backbone $G_{f}(\cdot;\theta_{f})$ away from the identity-related solution region and delivers a better solution for AU detection tasks.
@@ 5.2 F1 learning dynamics
In Sec. 3.2, we have mentioned the key design space of IAT: a linear projection layer for the ID head and a large $\lambda$ for the gradient reverse layer. These two factors together ensure the large magnitude of $-\lambda\frac{\partial L_{id}}{\partial\theta_{f}}$ during the adversarial training of the backbone $G_{f}(\cdot;\theta_{f})$. We elaborate here on the specifics of the IAT design space. We postulate that learning the subject identity is relatively easy, since there are many facial components and non-facial cues that can be used for identity recognition. Therefore, a strong regularization of IAT (i.e., a large $||-\lambda\frac{\partial L_{id}}{\partial\theta_{f}}||$) is required to counteract the identity-related learning tendency.
We first show the effect of using different $\lambda$ on the fold-2 of the BP4D dataset. All models share the same training settings except for $\lambda$. In Tab. 5, ‘Epoch’ indicates the training convergence point in terms of the F1 score and $\lambda=0$ represents the group without IAT. We see that a small $\lambda$ ($\lambda=0.02$), such as the one used in [87], has little gain of F1 score, whereas the large $\lambda$ ($\lambda=1,2,3$) yields significant improvement of AU prediction. Furthermore, the larger $\lambda$ we use, the more training epochs are required to reach a better optimization point, which is consistent with the phenomenon in Fig. 5.
@@
0.2 1 2 3
F1 68.33 68.60 68.63 69.26 69.57 69.47
Epoch 2 10 11 20 21 27
Furthermore, we show that recognizing identity is a trivial task since we find that a non-linear ID head $G_{f}(\cdot;\theta_{id})$ can still recognize the subjects given the identity-invariant features (regularized by IAT). To investigate this in more detail, we increase the model capacity of the ID head $G_{f}(\cdot;\theta_{id})$ given the backbone trained with a large $\lambda$, and measure the identity loss. Tab. 6 shows the results of using different MLP layers for FMAE-IAT under the same regularization strength ($\lambda=2$). The ID loss in Tab. 6 suggests that the model gradually learns the identity given some model capacity. By contrast, using the 1-layer MLP (linear projection layer) for the ID head leads to a large ID loss $L_{id}$, thus ensuring the large magnitude of $-\lambda\frac{\partial L_{id}}{\partial\theta_{f}}$. Therefore the linear projection layer is another necessity for IAT in AU detection. The convergence epoch and F1 score in Tab. 6 also imply that the 2-layer MLP and 3-layer MLP both converge fast and learn a sub-optimal solution to the AU tasks, which is consistent with our previous observations.
@@ 6 Conclusion
In conclusion, we have proposed to use a masked autoencoder (FMAE) with diverse pre-training for the AU detection task in this paper. We have leveraged a vast dataset (Face9M) for pretraining, combined with masked image modeling to significantly improve AU detection performance. Moreover, we have demonstrated the identity learning issue and its harmful effect in AU prediction models. The introduction of Identity Adversarial Training helped in mitigating the model’s learning of identity-related features. Also, We elucidated the two design factors of IAT, and our experiments consistently demonstrated superior performance over previous methods, achieving new SOTA results on AU benchmarks like BP4D, BP4D+ and DISFA.
We also noticed that the scaling effect of FMAE pretrained on Face9M has not converged even using the ViT-large model, pertaining by ViT-huge and distilling it into a smaller model for practical usage is promising, we leave it to future work.
We provide our trained models and our codebase as open-source to facilitate further research and application development. Note that some ethnicities or age groups may not be represented well in the large dataset. Users should exercise caution while using our pretrained models. Our models should not be used in applications to disadvantage minorities (e.g., develop systems to automatically hire employees by looking at their facial behavior during the interviews).

### v2

@@ 5 Analysis of Identity Adversarial Training
In this section, we elucidate IAT by first showing the identity learning issue in AU tasks (Section 5.1), then demonstrating the learning dynamics refined by IAT (Section 5.2), and finally illustrating the importance of ensuring a strong regularization of IAT (Section 5.3).
@@ 5.1 Linear probing for identity recognition
To demonstrate identity learning in AU detection, we quantitatively and qualitatively evaluate the identity features via linear probing [14] and t-SNE [67] technique, respectively. In detail, we apply linear probing on a trained AU detection model (FMAE) and evaluate the identity recognition accuracy on the BP4D dataset, which contains 41 subjects with the identity labels. Specifically, we freeze the backbone $G_{f}(\cdot;\theta_{f})$ of a well-trained FMAE and add a learnable linear classifier on top of the backbone to predict the identity label. For each subject in BP4D, we randomly draw 70 samples for training and 30 samples for testing. The resultant accuracy under linear probing is shown in Figure 3, the red line indicates that FMAE can recognize more than half of people among the 41 subjects even though the model is only trained for one epoch. Given enough training, the identity recognition accuracy can be as high as 83%. By contrast, IAT significantly alleviates this identity learning issue with 4.6% accuracy after one epoch of training and 27.9% accuracy at epoch 19. An interesting phenomenon is that even under the strong identity unlearning regularization, FMAE-IAT seems still to partially learn the identity-based features, by showing 27.9% accuracy (higher than the random guess accuracy 2.4%). We believe that the inherent high correlation between training and testing images for each subject provides the possibility for the model to infer the identity by looking at the non-face area.
Figure 3: Identity recognition accuracy (%) evaluated by linear probing on the BP4D dataset. IAT greatly reduces the identity-related features learned by the network backbone $G_{f}(\cdot;\theta_{f})$.
We also visualize the feature output from the backbone of FMAE and FMAE-IAT using t-SNE and see how these features are clustered according to the identity label. Figure 4 presents the t-SNE results for 20 subjects in BP4D (41 subjects in total), given trained FMAE and FMAE-IAT models. It is clear that the identity-based feature clusters in FMAE become less linearly distinguishable (the ID head is a linear layer) under the effect of IAT.
@@
5.2 IAT mitigates identity shortcut learning
After showing that a regular AU model (FMAE) learns the subject identity, we now illustrate that the identity shortcut learning leads to a trivial AU prediction solution that is inferior to the solution delivered by IAT. Concretely, we observe that FMAE and FMAE-IAT have totally different learning dynamics in terms of AU predictions (indicated by F1 score). Figure 5 shows the F1 score of both models along the training epochs, where the two models share the same learning rate, batch size and initial training states. It is clear from Figure 5 that FMAE is optimized quickly and converges at the third epoch with an F1 score of 65.45% under the identity shortcut. In contrast, FMAE-IAT learns the AU decision boundary progressively and converges only at epoch 15 with an F1 score of 66.66%. One can infer that IAT explicitly pushes the backbone $G_{f}(\cdot;\theta_{f})$ away from the identity-related solution region and delivers a better solution for AU detection tasks.
@@ 5.2 IAT mitigates identity shortcut learning
In Section 3.2, we have mentioned the key design space of IAT: a linear projection layer for the ID head and a large $\lambda$ for the gradient reverse layer. These two factors together ensure the large magnitude of $-\lambda\frac{\partial L_{id}}{\partial\theta_{f}}$ during the adversarial training of the backbone $G_{f}(\cdot;\theta_{f})$. We elaborate here on the specifics of the IAT design space. We postulate that learning the subject identity is relatively easy, since there are many facial components and non-facial cues that can be used for identity recognition. Therefore, a strong regularization of IAT (i.e., a large $||-\lambda\frac{\partial L_{id}}{\partial\theta_{f}}||$) is required to counteract the identity-related learning tendency.
We first show the effect of using different $\lambda$ on the fold-2 of the BP4D dataset. All models share the same training settings except for $\lambda$. In Table 5, ‘Epoch’ indicates the training convergence point in terms of the F1 score and $\lambda=0$ represents the group without IAT. We see that a small $\lambda$ ($\lambda=0.02$), such as the one used in [87], has little gain of F1 score, whereas the large $\lambda$ ($\lambda=1,2,3$) yields significant improvement of AU prediction. Moreover, the larger $\lambda$ we use, the more training epochs are required to reach a better optimization point, which is consistent with the phenomenon in Figure 5. Additionally, we perform the ablation study on $\lambda$ using the AU datasets to demonstrate that $\lambda$ is an easy hyper-parameter to tune in practice. Table 6 shows that $\lambda$ values within the set of [0.5, 1, 2], which are used across all AU datasets in this paper, consistently result in an improvement in the F1 score.
@@
1 2 3
F1 68.33 68.60 69.26 69.57 69.47
Epoch 2 10 20 21 27
Table 6: Ablation study of $\lambda$ on BP4D, BP4D+ and DISFA. F1 scores are reported for one fold of each dataset, with the numbers in parentheses indicating the absolute improvement in F1 compared to the baseline $\lambda=0$.
Dataset $\lambda$
0 0.5 1 2
BP4D 68.81 69.22 (+0.41) 69.26 (+0.45) 69.57 (+0.76)
BP4D+ 65.45 66.58 (+1.13) 66.66 (+1.21) 66.62 (+1.17)
DISFA 71.06 73.48 (+2.42) 73.23 (+2.17) 73.01 (+1.95)
Furthermore, we show that recognizing identity is a trivial task since we find that a non-linear ID head $G_{f}(\cdot;\theta_{id})$ can still recognize the subjects given the identity-invariant features (regularized by IAT). To investigate this in more detail, we increase the model capacity of the ID head $G_{f}(\cdot;\theta_{id})$ given the backbone trained with a large $\lambda$, and measure the identity loss. Table 7 shows the results of using different MLP layers for FMAE-IAT under the same regularization strength ($\lambda=2$). The ID loss in Table 7 suggests that the model gradually learns the identity given some model capacity. By contrast, using the 1-layer MLP (linear projection layer) for the ID head leads to a large ID loss $L_{id}$, thus ensuring the large magnitude of $-\lambda\frac{\partial L_{id}}{\partial\theta_{f}}$. Therefore the linear projection layer is another necessity for IAT in AU detection. The convergence epoch and F1 score in Table 7 also imply that the 2-layer MLP and 3-layer MLP both converge fast and learn a sub-optimal solution to the AU tasks, which is consistent with our previous observations.
@@ 6 Conclusion
In conclusion, we have proposed to use a masked autoencoder (FMAE) with diverse pre-training for the AU detection task in this paper. We have leveraged a vast and diverse dataset (Face9M) for pretraining, combined with masked image modeling to significantly improve AU detection performance. Moreover, we have demonstrated the identity learning issue and its harmful effect on AU prediction models. The use of Identity Adversarial Training helped in mitigating the model’s learning of identity-related features. Also, we elucidated the two design factors of IAT, and our experiments consistently demonstrated superior performance over previous methods, achieving new SOTA results on AU benchmarks like BP4D, BP4D+ and DISFA.
We also noticed that the scaling effect of FMAE pretrained on Face9M has not converged even using the ViT-large model. The potential of using ViT-huge and distilling it into a smaller model for practical use is promising, and we leave this for future work.